OpenStudy (anonymous):

Explain how the trinomial 2x2+ 13x + 15 can be turned into a four term polynomial suitable for factoring by grouping. Use complete sentences.

OpenStudy (mertsj):

Multiply the 2 and the 15 and find factors of 30 whose sum is 13 and then rewrite 13x as the sum of those two numbers. For example, 10 x3 = 30 and 10 + 3 is 13 so replace 13x with 10x + 3x

OpenStudy (anonymous):

2 * 15 = 30 - we need three factors which can be made to equal 13 - these are 2 , 3 and 5 ( 2*5 + 3) so we write 2x^2 + 3x + 5x + 15 we need 2*5 so we factor as follows 2x(x+5) + 3(x+5) = (2x + 3)(x + 5)

OpenStudy (mertsj):

Now you have 2x^2+10x+3x+15 Factor it by grouping